# Measurement of microwave photon correlations at millikelvin with a thermal detector

### v1

@@ Abstract
Microwave photons are important carriers of quantum information in many promising platforms for quantum computing. They can be routinely generated, controlled, and teleported in experiments, indicating a variety of applications in quantum technology. However, observation of quantum statistical properties of microwave photons remains demanding: The energy of several microwave photons is considerably smaller than the thermal fluctuation of any room-temperature detector, while amplification necessarily induces noise. Here, we present a measurement technique with a nanobolometer that directly measures the photon statistics at millikelvin and overcomes this trade-off. We apply our method to thermal states generated by a blackbody radiator operating in the regime of circuit quantum electrodynamics. We demonstrate the photon number resolvedness of the nanobolometer, and reveal the $n(n+1)$-scaling law of the photon number variance as indicated by the Bose–Einstein distribution. By engineering the coherent and incoherent proportions of the input field, we observe the transition between super-Poissonian and Poissonian statistics of the microwave photons from the bolometric second-order correlation measurement. This technique is poised to serve in fundamental tests of quantum mechanics with microwave photons and function as a scalable readout solution for a quantum information processor.
@@
Figure 1: Measurement scheme and characterization of the nanobolometer with coherent input. a Schematic diagram of the experimental setup. A beam splitter with a transmission rate $\Gamma=0.01$ combines a coherent microwave field at frequency $f_{\rm h}$ with a thermal field characterized by the radiation temperature $T$. The transmitted field is filtered and subsequently absorbed by a resistive absorber, causing a change of the electron temperature, $T_{\rm e}$, in the normal-metal nanowire. It leads to a decrease of the critical current of the SNS junctions (crosses) and hence a down-shift in the resonance frequency of the thermometer. This shift is measured by a weak probe field at frequency $f_{\rm p}$, which is coupled to the thermometer via a directional coupler (see Extend Data Fig. 1). b Reflection magnitude of the thermometer as a function of the probe power, $P_{\rm p}$, and the probe frequency, $f_{\rm p}$ (see Extend Data Fig. 2 for definition). The solid curve represents a cubic fit of the resonance frequency. c Reflection magnitude of the thermometer as a function of the input frequency, $f_{\rm h}$, and the probe frequency. The input and probe powers are fixed at $P_{\rm p}=-40\,{\rm dBm}$ and $P_{\rm h}=-20\,{\rm dBm}$, respectively. The solid curve is a Lorentzian fit of the resonance frequency, which indicates a passband of ${\rm FWHM}=133\,{\rm MHz}$ around the central frequency $f_{0}=8.428\,{\rm GHz}$. d Reflection magnitude of the thermometer as a function of the input power, $P_{\rm h}$, and the probe frequency, $f_{\rm p}$. Here, the input power is fixed at $P_{\rm p}=-40\,{\rm dBm}$ and $f_{\rm h}$ is set to the center frequency of the filter. The solid curve is a cubic fit of the resonance frequency.
@@
The nanobolometer, as shown in Fig. 1a, has a transmission-line filter that selects the bandwidth of interest. The filter is terminated by a normal-metal nanowire with an approximately $50\,{\Omega}$ resistance, i.e., the absorber. The absorber transforms the electromagnetic energy of the incident photons to a significant increase of the electron, or more precisely quasiparticle, temperature in the nanowire. This nanowire extends through several superconducting islands and forms an array of short SNS junctions that are embedded in a sub-gigahertz probe resonator, i.e., the thermometer [25]. The hot electrons weaken the coherent Andreev reflections at the SN and NS interfaces of each junction, and cause a down-shift of the resonance frequency of the thermometer. The amount of frequency shift is determined by the mean photon number of the input field, $\langle n\rangle$, considered in the unit of per second per hertz, which is the conventional unit for describing a stationary beam of electromagnetic field. The broadening of the line-shape is related to the photon number variance, $(\Delta n)^{2}=\langle n^{2}\rangle-\langle n\rangle^{2}$, similar to the Doppler broadening in laser spectroscopy [26] (see Methods). The nanobolometer therefore allows direct readout of the first two orders of photon number moments, $\langle n\rangle$ and $\langle n^{2}\rangle$, from the averaged scattering response of the thermometer without any amplification of the input field.
@@
Here, we use a cryogenic blackbody radiator to generate thermal microwave photons for bolometric measurement (see Methods). By controlling the local temperature of the radiator, $T$, we achieve an in-situ control of the mean radiation photon number according to the Planck’s law, $\langle n\rangle=1/\left\{\exp\left[hf_{\rm h}/(k_{\rm B}T)\right]-1\right\}$, where $f_{\rm h}$ is the input photon frequency, and $h$ and $k_{\rm B}$ are the Planck and Boltzmann constants, respectively. With the decrease of $\langle n\rangle$, the photon number variance converges from a quadratic function, $(\Delta n)^{2}\approx\langle n\rangle^{2}$ at $\langle n\rangle\gg 1$, to a linear relation, $(\Delta n)^{2}\approx\langle n\rangle$ at $\langle n\rangle\ll 1$[27]. Correspondingly, the photon statistics crosses over the boundary between the classical and the quantum regimes of the radiation field and converges to the shot-noise limit, indicating the discreteness of photons. This process is precisely described by the Bose-Einstein distribution of indistinguishable bosons, where $(\Delta n)^{2}=\langle n\rangle\left(\langle n\rangle+1\right)$.
@@
We also use a weak coherent input field as the reference of the bolometric measurement, where $(\Delta n)^{2}=\langle n\rangle$ by definition. Despite the different scaling laws of $(\Delta n)^{2}$, a stark contrast between the thermal and coherent photons lies in the second-order correlation function, $g^{(2)}(\tau)$. It characterizes the temporal separation between two successive photons in a propagating field. At zero delay, we have $g^{(2)}(0)=1+\left[(\Delta n)^{2}-\langle n\rangle\right]/\langle n\rangle^{2}$. A perfect coherent field obeys Poissonian statistics with $g^{(2)}(0)=1$, while thermal photons are super-Poissonian with $g^{(2)}(0)=2$. Considering that $g^{(2)}(\tau)=1$ for $\tau\rightarrow\infty$, an observation of $g^{(2)}(0)>1$ indicates the photon bunching effect where the photons tend to propagate in bundles.
@@
In our experiment, we first sweep the probe frequency from $f_{\rm p}=510$ to $530\,{\rm MHz}$ to identify the resonance frequency of the thermometer (Fig. 1b). At the minimum probe power, $P_{\rm p}=-60\,{\rm dBm}$ (see Extended Data Fig. 2 for definition), we extract the resonance frequency as $f_{\rm r}=524\,{\rm MHz}$, as well as the external and the total energy decay rates $\gamma_{\rm c}=4.8\,{\rm\mu s^{-1}}$ and $\gamma=18.7\,{\rm\mu s^{-1}}$, respectively, using the circle-fit method[28]. We observe that $f_{\rm r}$ remains almost invariant at $P_{\rm p}<-30\,{\rm dBm}$, but it starts to bend downwards at higher probe power together with the line-shape changing from a symmetric into an asymmetric form. Possible sources of this non-linearity are the coherent and incoherent Andreev reflections in the short SNS junctions, leading to the Josephson effect [29] and electrothermal feedback [30, 22], respectively.
We then set the probe power at $P_{\rm p}=-40\,{\rm dBm}$, and characterize the filter passband with a coherent input signal. Here, we sweep the input frequency from $f_{\rm h}=8.250$ to $8.600\,{\rm GHz}$ with a fixed power $P_{\rm h}=-20\,{\rm dBm}$ (Fig. 1c). Depending on the frequency of the input signal as compared to the passband, the resonance frequency of the thermometer, $f_{\rm r}$, shifts as a Lorentzian function of $f_{\rm h}$. The center frequency and bandwidth of the filter are determined to be $f_{0}=8.428\,{\rm GHz}$ and ${\rm FWHM}=133\,{\rm MHz}$, respectively.
@@
Figure 2: Photon-number resolved frequency shift. a Resonance frequency of the thermometer, $f_{\rm r}$, as a function of the probe power, $P_{\rm p}$, at different indicated temperatures of the blackbody radiation, $T$. The dots and solid curves are the corresponding experimental data and cubic fits, respectively. b Frequency shift, $\Delta f_{\rm r}$, in the region highlighted with grey in (a) as a function of the mean photon number of the thermal radiation field, $\langle n\rangle$. The dots and error bars represent the mean values and standard deviations of the raw data in the selected range. The dashed black line shows a cubic fit to the full data and the solid black line represents a linear fit to the data where $\langle n\rangle\geq 1$.
@@
For a more precise benchmark of the frequency shift with different $\langle n\rangle$, we use a blackbody radiator as the nanobolometer input. We sweep the radiation temperature, $T$, from the base temperature to approximately $1\,{\rm K}$ while keeping the mixing chamber (MXC) temperature below $60\,{\rm mK}$. The corresponding input photon number, $\langle n\rangle$, can be precisely determined by the Planck’s law, which varies from 00 to $2\,{\rm photon/(s\times Hz)}$. At each radiation temperature, we sweep both $P_{\rm p}$ and $f_{\rm p}$ and identify the resonance frequency, $f_{\rm r}$, at the dip of the reflection magnitude, $|S_{11}|$. The intermediate-frequency (IF) bandwidth is set to $200\,{\rm Hz}$ with $10$ averages. Figure 2a summarizes the measurement results. A clear decrease of $f_{\rm r}$ with increasing $T$ is observed, which is consistent with the observation with a coherent input field as shown in Fig. 1d.
@@
Averaging the relative frequency shift within a ±$\pm 3\,{\rm dB}$ range around $P_{\rm p}=-40\,{\rm dBm}$, we observe a cubic relation between $\Delta f_{\rm r}$ and $\langle n\rangle$ (Fig. 2b). At the minimum value of $\langle n\rangle=0.16\,{\rm photon/(s\times Hz)}$, we obtain the mean and standard deviation of $-\Delta f_{\rm r}$ as $0.22\,{\rm MHz}$ and $0.19\,{\rm MHz}$, respectively, corresponding to a coefficient of variation (CV) less than $1$. We therefore attribute $0.16\,{\rm photon/(s\times Hz)}$ as an upper bound of the photon number resolvedness of the nanobolometer. The value of CV decreases monotonically with $\langle n\rangle$ and approaches $0.05$ at the highest photon number. We note that the relatively small photon number resolution is benefit from the broad bandwidth of the blackbody radiation that covers the entire $133\,{\rm MHz}$ passband of the input filter. The corresponding input-field-independent power resolution of the nanobolometer is $119\,{\rm aW}$. In fact, this resolution may be pushed to an even smaller value by decreasing the IF bandwidth or increasing the number of averages.
@@
Figure 3: Photon number variance extracted from the reflection spectrum. a Photon number variance, $(\Delta n)^{2}$, as a function of the mean input photon number, $\langle n\rangle$, for the thermal (circles) and coherent (triangles) input fields. The solid and dashed lines represent the theoretical expectation for thermal and coherent states, respectively. The thermal photon number is obtained by Planck’s law, and the coherent photon number is obtained by assuming an $88\,{\rm dB}$ attenuation in the input line as a scaling parameter. The detection efficiency of the nanobolometer is assumed to be unity to minimize the number of fitting parameters. The inset shows the region near the origin. b Observed frequency shift, $\Delta f_{\rm r}$, as a function of $\langle n\rangle$ for thermal (circles) and coherent (triangles) input fields. The open circles are corrected results by considering the phonon temperature change, $T_{\rm b}$, during the temperature sweep. The inset shows the records of two different sensors that are mounted at the sample holder and the MXC, respectively. The former is lower-bounded by approximately $45\,{\rm mK}$ due to self-heating.
@@
The quantum statistical properties of thermal states are fully determined by the mean photon number, $\langle n\rangle$, and its variance, $(\Delta n)^{2}$. With a weak radio-frequency (RF) probe field applied to the thermometer, these two values can be respectively obtained from the fitted resonance frequency, $\mu$, and the broadening of the spectrum linewidth, $\sigma^{2}$ (see Methods). Here, we sweep the radiation temperature, $T$, from the base temperature to approximately $2\,{\rm K}$ and measure the averaged reflection coefficient, $S_{11}$, at a fixed probe power, $P_{\rm p}=-40\,{\rm dBm}$. We first fit the theoretical curve to the data at the minimum $T$ as a reference curve for zero broadening ($\sigma^{2}=0$). We then extract $\sigma^{2}$ at different values of $T$ by comparing the measurement results with this reference (see Methods). The photon number variance is obtained by assuming a linear relation $\Delta n=\alpha\sigma$, where $\alpha=1.92\,{\rm photon/MHz}$ is a fitted scaling factor (see Methods).
@@
At a sufficiently small photon number, the extracted photon number variance, $(\Delta n)^{2}$, is almost equal to $\langle n\rangle$, as shown in the inset of Fig. 3a. It reveals the particle nature of photons in the form of shot noise, as described by the Poisson distribution. On the other hand, $(\Delta n)^{2}$ can be well approximated by a quadratic function, $\langle n\rangle^{2}$, at large $\langle n\rangle$. These two limits are connected by the Bose–Einstein distribution, which predicts $(\Delta n)^{2}\approx\langle n\rangle\left(\langle n\rangle+1\right)$ for indistinguishable bosons in thermal states. Our experimental results in Fig. 3a show an excellent agreement with the theoretical expectation. It reveals the transition between the quantum and classical regimes of a microwave field with the increase of the photon number.
@@
For comparison, we replace the thermal input field by a coherent signal at $f_{0}$. We sweep the input power from $P_{\rm h}=-40$ to $-20\,{\rm dBm}$, and extract $\langle n\rangle$ and $(\Delta n)^{2}$ with almost identical fitting parameters as for the thermal fields (see Extended Data Fig. 3 for comparison). We observe a qualitatively different statistics of photons between the thermal and coherent input fields, as shown in Fig. 3a. Here, $(\Delta n)^{2}$ remains as a pure linear function of $\langle n\rangle$ in the entire range of the power sweep, corresponding to a mean photon number from 00 to $19$. This observation is in agreement with the Poisson statistics of the input photons in a coherent state. The excellent agreement between theory and experiments for both thermal and coherent states with almost the same fitting parameter demonstrate the correctness of our measurement result.
@@
Figure. 3b shows the extracted value of $\Delta f_{\rm r}$ for both thermal and coherent inputs. The data are in quantitative agreement with each other for $\langle n\rangle<2$, indicating an almost linear relation between the frequency shift of the thermometer and the input power as observed in Fig. 2b. However, a noticeable discrepancy in $\Delta f_{\rm r}$ is observed at the higher input power. Since the temperature of the blackbody radiator, $T$, is significantly higher than the MXC temperature in this range, we attribute the observed discrepancy to the heating of the phonon bath temperature, $T_{\rm b}$. The latter is recorded in the inset of Fig. 3b. We consider a linear contribution of $T_{\rm b}$ to the net heating power of the nanobolometer, i.e., $P=\beta T_{\rm b}+{\rm FWHM}k_{\rm B}T$. We obtain a quantitative agreement of $\Delta f_{\rm r}$ for the two types of photons with $\beta=440\,{\rm fW/K}$, as shown in Fig. 3b.
@@
Figure 4: Second-order correlation for different photon states. The proportions of the coherent and thermal parts in the input field are controlled by varying the power of the coherent field, $P_{\rm h}$, and the radiation temperature, $T$. Here, the solid curves correspond to the theoretical expectations at fixed values of $P_{\rm h}$ but with a varying $T$. The inset shows the raw data of the photon number and variance that are used to calculate $g^{(2)}(0)$.
@@
The shown ability of resolving the mean photon number, $\langle n\rangle$, and its variance, $(\Delta n)^{2}$, indicates a convenient way of measuring the zero-delay second-order correlation functions of the incident microwave field, $g^{(2)}(0)$. In our experiment, we control the coherent and incoherent proportions of the input field by adjusting their powers at the beam splitter inputs (Fig. 1a), and repeat the photon variance measurement as shown in Fig. 3.

### v2

# Correlation measurement of propagating microwave photons at millikelvin

@@ Abstract
Microwave photons are important carriers of quantum information in many promising platforms for quantum computing. They can be routinely generated, controlled, and teleported in experiments, indicating a variety of applications in quantum technology. However, observation of quantum statistical properties of microwave photons remains demanding: The energy of several microwave photons is considerably smaller than the thermal fluctuation of any room-temperature detector, while amplification necessarily induces noise. Here, we present a measurement technique with a nanobolometer that directly measures the photon statistics at the millikelvin temperature and overcomes this trade-off. We apply our method to thermal states generated by a blackbody radiator operating in the regime of circuit quantum electrodynamics. We demonstrate the photon number resolvedness of the nanobolometer, and reveal the $n(n+1)$-scaling law of the photon number variance as indicated by the Bose–Einstein distribution. By engineering the coherent and incoherent proportions of the input field, we observe the transition between super-Poissonian and Poissonian statistics of the microwave photons from the bolometric second-order correlation measurement. This technique is poised to serve in fundamental tests of quantum mechanics with microwave photons and function as a scalable readout solution for a quantum information processor.
@@
Figure 1: Measurement scheme and characterization of the nanobolometer with coherent input. a Schematic diagram of the experimental setup. A beam splitter with a transmission rate $\Gamma=0.01$ combines a coherent microwave field at frequency $f_{\rm h}$ with a thermal field characterized by the radiation temperature $T$. The transmitted field is filtered and subsequently absorbed by a resistive absorber, causing a change of the electron temperature, $T_{\rm e}$, in the normal-metal nanowire. It leads to a decrease of the critical current of the SNS junctions (crosses) and hence a down-shift in the resonance frequency of the thermometer. This shift is measured by a weak probe field at frequency $f_{\rm p}$, which is coupled to the thermometer via a directional coupler (see Extend Data Fig. 1). b Reflection magnitude of the thermometer as a function of the probe power at the sample, $P_{\rm p}$, and the probe frequency, $f_{\rm p}$. The solid curve represents a cubic fit of the resonance frequency. c Reflection magnitude of the thermometer as a function of the input frequency, $f_{\rm h}$, and the probe frequency. The probe and input powers are fixed at $P_{\rm p}=-120\,{\rm dBm}$ and $P_{\rm h}=-108\,{\rm dBm}$, respectively. The solid curve is a Lorentzian fit of the resonance frequency, which indicates a passband of ${\rm FWHM}=133\,{\rm MHz}$ around the central frequency $f_{0}=8.428\,{\rm GHz}$. d Reflection magnitude of the thermometer as a function of the input power at the sample, $P_{\rm h}$, and the probe frequency, $f_{\rm p}$. Here, the probe power is fixed at $P_{\rm p}=-120\,{\rm dBm}$ and $f_{\rm h}$ is set to the center frequency of the filter. The solid curve is a cubic fit of the resonance frequency. The data is normalized to $0\,{\rm dB}$ at $530\,{\rm MHz}$ probe frequency in (b)–(d).
@@
The nanobolometer, as shown in Fig. 1a, has a transmission-line filter that selects the bandwidth of interest. The filter is terminated by a normal-metal nanowire with an approximately $50\,{\Omega}$ resistance, i.e., the absorber. The absorber transforms the electromagnetic energy of the incident photons to a significant increase of the electron temperature in the nanowire. This nanowire extends through several superconducting islands and forms an array of short SNS junctions that are embedded in a sub-gigahertz probe resonator, i.e., the thermometer [25]. The hot electrons weaken the coherent Andreev reflections at the SN and NS interfaces of each junction, and cause a down-shift of the mean resonance frequency of the thermometer. The amount of the shift is determined by the mean absorbed power, which for a given frequency and bandwidth can be characterized by the mean photon number of the input field, $\langle\hat{n}\rangle$, considered in the unit of per second per hertz as it is the conventional unit for describing a stationary beam of electromagnetic field (see Supplementary Note 1). The broadening of the line-shape is related to the photon number variance, $(\Delta n)^{2}=\langle\hat{n}^{2}\rangle-\langle\hat{n}\rangle^{2}$, similar to the Doppler broadening in laser spectroscopy [26] (see Methods). The nanobolometer therefore allows direct readout of the first two orders of photon number moments, $\langle\hat{n}\rangle$ and $\langle\hat{n}^{2}\rangle$, from the long-time-averaged scattering response of the thermometer without any amplification of the input field. Note that the incident and probe fields are not directly electrically coupled to each other, owing to the ground port between the absorber and the thermometer, and have several gigahertz difference in frequency.
@@
Here, we use a cryogenic blackbody radiator to generate thermal microwave photons for bolometric measurement (see Methods). By controlling the local temperature of the radiator, $T$, we achieve an in-situ control of the mean radiation photon number according to the Planck’s law, $\langle\hat{n}\rangle=1/\left\{\exp\left[hf_{\rm h}/(k_{\rm B}T)\right]-1\right\}$, where $f_{\rm h}$ is the input photon frequency, and $h$ and $k_{\rm B}$ are the Planck and Boltzmann constants, respectively. With the decrease of $\langle\hat{n}\rangle$, the photon number variance converges from a quadratic function, $(\Delta n)^{2}\approx\langle\hat{n}\rangle^{2}$ at $\langle\hat{n}\rangle\gg 1$, to a linear relation, $(\Delta n)^{2}\approx\langle\hat{n}\rangle$ at $\langle\hat{n}\rangle\ll 1$[27]. Correspondingly, the photon statistics crosses over the boundary between the classical and the quantum regimes of the radiation field and converges to the shot-noise limit, indicating the discreteness of photons. This process is precisely described by the Bose-Einstein distribution of indistinguishable bosons, where $(\Delta n)^{2}=\langle\hat{n}\rangle\left(\langle\hat{n}\rangle+1\right)$.
@@
We also use a weak coherent input field as the reference of the bolometric measurement, where $(\Delta n)^{2}=\langle\hat{n}\rangle$ by definition. Despite the different scaling laws of $(\Delta n)^{2}$, a stark contrast between the thermal and coherent photons lies in the second-order correlation function, $g^{(2)}(\tau)$, with $\tau$ being the time delay. It characterizes the temporal separation between two successive photons in a propagating field. At zero delay, we have $g^{(2)}(0)=1+\left[(\Delta n)^{2}-\langle\hat{n}\rangle\right]/\langle\hat{n}\rangle^{2}$. A perfect coherent field obeys Poissonian statistics with $g^{(2)}(0)=1$, while thermal photons are super-Poissonian with $g^{(2)}(0)=2$. Considering that $g^{(2)}(\tau)=1$ for $\tau\rightarrow\infty$, an observation of $g^{(2)}(0)>1$ indicates the photon bunching effect where the photons tend to propagate in bundles.
Results
@@
In our experiment, we first sweep the probe frequency from $f_{\rm p}=510$ to $530\,{\rm MHz}$ to identify the resonance frequency of the thermometer (Fig. 1b). At the minimum probe power, $P_{\rm p}=-140\,{\rm dBm}$ defined at the thermometer input by assuming an $80\,{\rm dB}$ attenuation from the source, we extract the resonance frequency as $f_{\rm r}=524\,{\rm MHz}$, as well as the external and the total energy decay rates $\gamma_{\rm c}=4.8\,{\rm\mu s^{-1}}$ and $\gamma=18.7\,{\rm\mu s^{-1}}$, respectively, using the circle-fit method[28]. We observe that $f_{\rm r}$ remains almost invariant at $P_{\rm p}<-110\,{\rm dBm}$, but it starts to bend downwards at higher probe power together with the line-shape changing from a symmetric into an asymmetric form. Possible sources of this non-linearity are the coherent and incoherent Andreev reflections in the short SNS junctions, leading to the Josephson effect [29] and electrothermal feedback [30, 22], respectively.
We then set the probe power at $P_{\rm p}=-120\,{\rm dBm}$, and characterize the filter passband with a coherent input signal. Here, we sweep the input frequency from $f_{\rm h}=8.25$ to $8.60\,{\rm GHz}$ with a fixed power $P_{\rm h}=-108\,{\rm dBm}$ defined at the filter input by fitting an $88\,{\rm dB}$ attenuation from the source (Fig. 1c). Depending on the frequency of the input signal as compared to the passband, the resonance frequency of the thermometer, $f_{\rm r}$, shifts as a Lorentzian function of $f_{\rm h}$. The center frequency and bandwidth of the filter are determined to be $f_{0}=8.428\,{\rm GHz}$ and ${\rm FWHM}=133\,{\rm MHz}$, respectively.
@@
Figure 2: Photon-number resolved frequency shift. a Resonance frequency of the thermometer, $f_{\rm r}$, as a function of the probe power, $P_{\rm p}$, at different indicated temperatures of the blackbody radiation, $T$. The dots and solid curves are the corresponding experimental data and cubic fits, respectively. b Frequency shift, $\Delta f_{\rm r}$, in the region highlighted with grey in (a) as a function of the mean photon number of the thermal radiation field, $\langle\hat{n}\rangle$. The dots and error bars represent the mean values and standard deviations of the raw data in the selected range. The dashed black line shows a cubic fit to the full data and the solid black line represents a linear fit to the data where $\langle\hat{n}\rangle\geq 1$. The top horizontal axis indicates the radiation power in the entire $133\,{\rm MHz}$ bandwidth of the input filter.
@@
For a more precise benchmark of the frequency shift with different $\langle\hat{n}\rangle$, we use a blackbody radiator as the nanobolometer input. We sweep the radiation temperature, $T$, from the base temperature to approximately $1\,{\rm K}$ while keeping the mixing chamber (MXC) temperature below $60\,{\rm mK}$. The corresponding input photon number, $\langle\hat{n}\rangle$, can be precisely determined by the Planck’s law, which varies from 00 to $2\,{\rm photons/(s\times Hz)}$. At each radiation temperature, we sweep both $P_{\rm p}$ and $f_{\rm p}$ and identify the resonance frequency, $f_{\rm r}$, at the dip of the averaged reflection magnitude. Here, the intermediate-frequency (IF) bandwidth is set to $200\,{\rm Hz}$ with $10$ averages. Figure 2a summarizes the measurement results. A clear decrease of $f_{\rm r}$ with increasing $T$ is observed, which is consistent with the observation with a coherent input field as shown in Fig. 1d.
@@
Averaging the relative frequency shift within a ±$\pm 3\,{\rm dB}$ range around $P_{\rm p}=-120\,{\rm dBm}$, we observe a cubic relation between $\Delta f_{\rm r}$ and $\langle\hat{n}\rangle$ (Fig. 2b). At the minimum value of $\langle\hat{n}\rangle=0.16\,{\rm photons/(s\times Hz)}$, we obtain the mean and standard deviation of $-\Delta f_{\rm r}$ as $0.22\,{\rm MHz}$ and $0.19\,{\rm MHz}$, respectively, corresponding to a coefficient of variation (CV) less than $1$. We therefore attribute $0.16\,{\rm photons/(s\times Hz)}$ as an upper bound of the photon number resolvedness of the nanobolometer. The value of CV decreases monotonically with $\langle\hat{n}\rangle$ and approaches $0.05$ at the highest photon number. In fact, this resolution may be pushed to an even smaller value by decreasing the IF bandwidth or increasing the number of averages.
We note that the relatively small photon number resolution is benefited from the broad bandwidth of the blackbody radiation that covers the entire $133\,{\rm MHz}$ passband of the input filter. The corresponding input-field-independent power resolution of the nanobolometer is $119\,{\rm aW}$, corresponding to a considerable amount of individual photon wavepackets per second, $2.1\times 10^{7}\,{\rm photons/s}$. In this regard, the nanobolometer reported here is still far from resolving each individual photon emission event in a microwave field, which has been demonstrated in the optical regime  [31]. Fortunately, the observed power resolution is adequate for resolving the quantum statistical properties of propagating microwave photons that are conventionally defined by the arriving photons per second per hertz [27] (see Supplementary Note 1).
@@
Figure 3: Photon number variance extracted from the reflection spectrum. a Photon number variance, $(\Delta n)^{2}$, as a function of the mean input photon number, $\langle\hat{n}\rangle$, for the thermal (circles) and coherent (triangles) input fields. The solid and dashed lines represent the theoretical expectation for thermal and coherent states, respectively. The thermal photon number is obtained by Planck’s law, and the coherent photon number is obtained by assuming an $88\,{\rm dB}$ attenuation in the input line as a scaling parameter. The detection efficiency of the nanobolometer is assumed to be unity to minimize the number of fitting parameters. The inset shows the region near the origin. b Extracted resonance frequency, $\mu$, as a function of $\langle\hat{n}\rangle$ for thermal (circles) and coherent (triangles) input fields. The open circles are corrected results by considering the phonon temperature change, $T_{\rm b}$, during the temperature sweep. The inset shows the records of two different sensors that are mounted at the sample holder and the MXC, respectively. The former is lower-bounded by approximately $45\,{\rm mK}$ due to self-heating. The top horizontal axis indicates the radiation power in the entire $133\,{\rm MHz}$ bandwidth of the input filter.
@@
The quantum statistical properties of thermal states are fully determined by the mean photon number, $\langle\hat{n}\rangle$, and its variance, $(\Delta n)^{2}$. With a weak radio-frequency (RF) probe field applied to the thermometer, these two values can be respectively obtained from the fitted resonance frequency, $\mu$, and the broadening of the spectrum linewidth, $\sigma^{2}$ (see Methods). Here, we sweep the radiation temperature, $T$, from the base temperature to approximately $2\,{\rm K}$ and measure the averaged reflection coefficient, $S_{11}^{\rm ave}$, at a fixed probe power, $P_{\rm p}=-120\,{\rm dBm}$. We first fit the theoretical curve to the data at the minimum $T$ as a reference curve for zero broadening ($\sigma^{2}=0$). We then allow $\sigma^{2}$ to be a fitting parameter and extract its value at different $T$ by comparing the measurement results with this reference (see Supplementary Note 4). The photon number variance is obtained by assuming a linear relation $\Delta n=\alpha\sigma$, where $\alpha=1.97\,{\rm photons/MHz}$ is a fitted scaling factor.
@@
At a sufficiently small photon number, where $\langle\hat{n}\rangle$ is precisely determined by the Planck’s law, the extracted photon number variance, $(\Delta n)^{2}$, is almost equal to $\langle\hat{n}\rangle$, as shown in the inset of Fig. 3a. It reveals the particle nature of photons in the form of shot noise, as described by the Poisson distribution. On the other hand, $(\Delta n)^{2}$ can be well approximated by a quadratic function, $\langle\hat{n}\rangle^{2}$, at large $\langle\hat{n}\rangle$. These two limits are connected by the Bose-Einstein distribution, which predicts $(\Delta n)^{2}\approx\langle\hat{n}\rangle\left(\langle\hat{n}\rangle+1\right)$ for indistinguishable bosons in thermal states. Our experimental results in Fig. 3a show an excellent agreement with the theoretical expectation. It reveals the transition between the quantum and classical regimes of a microwave field with the increase of the photon number.
@@
For comparison, we replace the thermal input field by a coherent signal at $f_{0}$. We sweep the input power from $P_{\rm h}=-128$ to $-108\,{\rm dBm}$, and extract $\langle\hat{n}\rangle$ and $(\Delta n)^{2}$ with almost identical fitting parameters as for the thermal fields (see Supplementary Note 4). We observe a qualitatively different statistics of photons between the thermal and coherent input fields, as shown in Fig. 3a. Here, $(\Delta n)^{2}$ remains as a pure linear function of $\langle\hat{n}\rangle$ in the entire range of the power sweep, corresponding to a mean photon number from 00 to $19$. This observation is in agreement with the Poisson statistics of the input photons in a coherent state. The excellent agreement between theory and experiments for both thermal and coherent states with almost the same fitting parameter demonstrate the correctness of our measurement result.
@@
Figure 3b shows the extracted value of $\mu$ for both thermal and coherent inputs. The data are in quantitative agreement with each other for $\langle\hat{n}\rangle<2$, indicating an almost linear relation between the frequency shift of the thermometer and the input power as observed in Fig. 2b. However, a noticeable discrepancy is observed at the higher input power. Since the temperature of the blackbody radiator, $T$, is significantly higher than the MXC temperature in this range, we attribute the observed discrepancy to the heating of the phonon bath temperature, $T_{\rm b}$. The latter is recorded in the inset of Fig. 3b. We then consider a $6^{\rm th}$-order polynomial correction of the net heating power due to the phonon temperature change, i.e., $P=\beta T_{\rm b}^{6}+P_{\rm h}$, for ultra-thin films [32]. A qualitative agreement of $\mu$ for the two types of photons is achieved with $\beta=1.84\,{\rm nW/K^{6}}$, as shown in Fig. 3b. A quantitative agreement can be achieved by assuming an approximately $50\%$ absorption rate of the nanobolometer, which is not taken into account throughout this study.
@@
Figure 4: Second-order correlation for different photon states. The proportions of the coherent and thermal parts in the input field are controlled by varying the power of the coherent field, $P_{\rm h}$, and the radiation temperature, $T$. Here, the solid curves correspond to the theoretical expectations at fixed values of $P_{\rm h}$ but with a varying $T$. The inset shows the raw data of the photon number and variance that are used to calculate $g^{(2)}(0)$. The top horizontal axis indicates the radiation power in the entire $133\,{\rm MHz}$ bandwidth of the input filter.
@@
The shown ability of resolving the mean photon number, $\langle\hat{n}\rangle$, and its variance, $(\Delta n)^{2}$, indicates a convenient way of measuring the zero-delay second-order correlation functions of the incident microwave field, $g^{(2)}(0)$. In our experiment, we control the coherent and incoherent proportions of the input field by adjusting their powers at the beam splitter inputs (Fig. 1a), and repeat the photon variance measurement as shown in Fig. 3.